Find the sum of 395 & 276 how do you do that please help
1
Expert's answer
2012-08-28T08:42:08-0400
Obviously, we could use calculator for that. But if you don't have one or are not allowed to use it in this task or whatever, you can use long addition. let's see

395
& +
& 276
__________
& 671

You write your numbers one above another so that their last (right) digits were strictly one above other. And you begin with the last digits adding upper and lover one. In our case we firstly add 5+6=11; we write 1 under the line and another 1 we remember. Then we add 9+7=16 and also add that 1 we remembered. so we get 16+1=17. we write 7 in the answer field under the line, 1 again is remembered. Finally, we add 3+2=5 and also 1 so we get 5+1=6. we write 6 in the answer field. So we get 671